# LayeredMAPF: a decomposition of MAPF instance without compromising solvability

### v1

@@ V-C Time cost and memory usage
Generally, as depicted in the Fig. 3 and Fig. 4, the time cost of decomposition is less than 1 second in most maps and less than 3 seconds in the worst-case scenario (such as 800 to 1000 dense agents), which is deemed acceptable for practical application. In the majority of cases, all three steps contribute to the time cost. However, there are differences in the time cost of steps across different maps. In maps with an abundance of free grids exceeding the number of agents, such as den512d, Berlin_1_256, and Paris_1_256, step 1 consumes almost all of the time cost, resulting in overlapping time costs after step 1, step 2, and step 3. Conversely, in maps with few free grids close to the number of agents, all steps contribute to the time cost, as these scenarios require steps 2 and 3 to decompose initial subproblems into smaller subproblems.
@@
V-D Summary
@@
Our method demonstrates high effectiveness for MAPF instances with free grids exceeding two times the number of agents, while its effectiveness diminishes as the number of free grids approaches two times the number of agents. This is because a higher number of free grids increases the likelihood of decomposing the instance into smaller subproblems. Conversely, for the same map, the effectiveness of decomposition decreases with an increase in the number of agents in the MAPF instance.
@@
In terms of costs, the average time cost is less than 1 second in average, with a maximum of less than 3 seconds in the worst cases (such as 800 to 1000 dense agents), and memory usage remains below 1 MB. Comparatively, maps with more free grids than the number of agents require fewer computations and less memory space.
VI Results of decomposition’s application
[…]
Figure 5: These figures provide a summary comparison between the raw MAPF method and the layered MAPF method, combining the results across all maps. In the visualizations, crossed bars represent the layered version of the MAPF method, while empty bars represent the raw MAPF method. The same applies hereinafter.
@@ VI Results of decomposition’s application
VI-A Explicit Estimation CBS
Explicit Estimation CBS (EECBS) [li2021eecbs] stands out as one of the state-of-the-art CBS-based MAPF methods, characterized by its completeness and bounded suboptimality. Beyond its explicit estimation in high-level search, EECBS also incorporates conflict resolution techniques such as bypassing conflicts, prioritizing conflicts, symmetry reasoning, and weighted dependency graphs to reduce the time cost of CBS. Further details can be found in Section II. EECBS’s code  [^3] https://github.com/Jiaoyang-Li/EECBS operates by searching each agent’s path separately and provides interfaces to avoid conflicts with external paths, rendering it a serial MAPF method. In its layered version, which we refer to as ”layered EECBS,” previous subproblems’ paths and subsequent subproblems’ starts are set as constraints.
As illustrated in Fig. 5, on average, layered EECBS exhibits lower time costs ($1.0\times 10^{4}$ ms $<$ $1.5\times 10^{4}$ ms), leading to a higher success rate (0.77 $>$ 0.64). The time cost of decomposition is relatively small compared to the time cost of solving the MAPF instance. Regarding memory usage, layered EECBS consumes less memory space compared to raw EECBS ($12$ MB $<$ $24$ MB). Decomposition of the MAPF instance significantly reduces the growth of the conflict tree’s size by decreasing the number of agents solved simultaneously.
In terms of path quality, layered EECBS yields a larger makespan (layered EECBS: $2.5\times 10^{2}$, raw EECBS: $2.0\times 10^{2}$), resulting in a larger sum of costs (layered EECBS: $5.0\times 10^{4}$, raw EECBS: $3.4\times 10^{4}$). In maps where both raw EECBS and layered EECBS achieve a success rate close to 1 (e.g., map 1, 8, 10, 19, 21 and 22), both methods exhibit similar makespan and sum of costs. However, in maps where raw EECBS has a lower success rate compared to layered EECBS (e.g., map 4, 16, 17 and 24), layered EECBS demonstrates a larger makespan and sum of costs, as depicted in Fig. 6.
@@ VI-A Explicit Estimation CBS
VI-B Priority Based Search
Priority Based Search (PBS) [ma2019searching] is an incomplete, suboptimal, and priority-based MAPF method. Conceptually, when faced with a collision, PBS greedily chooses which agent should be given higher priority. It backtracks efficiently and explores other branches only if there is no solution in the current branch. PBS searches each agent’s path separately and it’s code [^4] https://github.com/Jiaoyang-Li/PBS provides interfaces to avoid conflicts with external paths, making it a serial MAPF method. In its layered version, which we refer to as ”layered PBS,” we set previous subproblems’ paths and subsequent subproblems’ starts as constraints.
As shown in Fig. 5, on average, layered PBS has a lower time cost ($1.2\times 10^{4}$ ms $<$ $2.0\times 10^{4}$ ms) and thus a higher success rate (0.75 $>$ 0.42). The time cost of decomposition is relatively small compared to the time cost of solving the MAPF instance. In terms of memory usage, layered PBS has lower memory usage compared to raw PBS (1.0 MB $<$ 3.9 MB) because decomposition of the MAPF instance reduces the complexity of the priority tree significantly.
When it comes to path quality, layered PBS has a larger makespan ($2.5\times 10^{2}$ $>$ $1.3\times 10^{2}$) and thus a larger SOC ($5.5\times 10^{4}$ $>$ $1.7\times 10^{4}$). Similar to EECBS, raw PBS finds it difficult to find longer solutions and can only finds shorter solutions, resulting in lower success rates for maps where raw PBS struggles (e.g., map 4, 6, 16 and 17). For maps where both methods have high success rates (e.g., map 1, 9 and 10), they have similar makespan and SOC, as shown in Fig. 6.
@@ VI-B Priority Based Search
VI-C Large Neighborhood Search 2
Large Neighborhood Search 2 (LNS2) [li2022mapf] is a complete and suboptimal MAPF method that starts from a set of paths with collisions and repeatedly replans subsets of paths to reduce the overall number of collisions until the paths become collision-free. LNS2 updates each agent’s path separately, and its code [^5] https://github.com/Jiaoyang-Li/MAPF-LNS provides interfaces to avoid conflicts with external paths, making it a serial MAPF method.
As shown in Fig. 5, on average, layered LNS2 has a lower time cost ($1.0\times 10^{4}$ ms $<$ $1.4\times 10^{4}$ ms) and thus a higher success rate (0.76 $>$ 0.55). In terms of memory usage, layered LNS2 has lower memory usage compared to raw LNS2 (2.4 MB $<$ 32 MB). Due to the decomposition of the MAPF instance, other agents’ paths are considered as dynamic obstacles and do not join conflict resolve, resulting in layered LNS2 requiring fewer repair actions to fix conflicts between agents’ paths.
Regarding path quality, layered LNS2 has a larger makespan ($2.5\times 10^{2}$ $>$ $2.1\times 10^{2}$) and SOC ($5.1\times 10^{4}$ $>$ $3.4\times 10^{4}$). Similar to EECBS and PBS, raw LNS2 struggles to find longer solutions and can only discovers shorter solutions, particularly in maps where raw LNS2 has a lower success rate (e.g., map 12 and 16). For maps where both methods have similar success rates (e.g., map 1, 2, 8, 21 and 22), raw LNS2 and layered LNS2 have similar makespan and SOC, as shown in Fig. 6.
@@ VI-C Large Neighborhood Search 2
VI-D Push and Swap
Push and Swap (PAS) [luna2011push] is a suboptimal and complete rule-based MAPF method. The algorithm employs two primitives: a “push” operation where agents move towards their goals until no progress can be made, and a “swap” operation that allows two agents to exchange positions without affecting other agents’ configurations.
For each agent in the MAPF instance, the algorithm uses “push” to move the agent toward its goal along the shortest path until encountering another agent. Then, it employs “swap” to exchange their positions without altering other agents’ positions. If a “swap” fails, the problem is deemed unsolvable. The algorithm alternates between ”push” and “swap” operations until all agents reach their targets.
Push and Swap updates agents’ paths simultaneously (via an action called “MULTIPUSH”, which moves a set of adjacent agents simultaneously), and its code [^6] https://github.com/Kei18/pibt2 provides no interfaces to avoid conflicts with external paths and ensures that the same state of agents does not occur twice in the solution, making it a parallel MAPF method. In its layered version, referred to as ”layered Push and Swap” or ”layered PAS,” we set previous subproblems’ targets and subsequent subproblems’ starts as static obstacles. Subsequently, we merge subproblems’ solutions by adding wait actions after solving all subproblems.
As shown in Fig. 5, on average, layered PAS consumes less time ($4.6\times 10^{3}$ ms $<$ $8.8\times 10^{3}$ ms) and memory space (6.2 MB $<$ 70 MB) than raw PAS, resulting in a higher success rate (0.95 $>$ 0.82). Decomposing the MAPF instance enables layered Push and Swap to solve subproblems separately, requiring fewer “push” and “swap” actions. However, the number of “push” and “swap” actions grows exponentially as the number of agents increases.
Regarding path quality, as PAS is a parallel MAPF method, the addition of wait actions to merge subproblems’ solutions, layered PAS has a larger makespan ($1.0\times 10^{4}$ $>$ $6.1\times 10^{3}$) and SOC ($3.6\times 10^{6}$ $>$ $2.4\times 10^{6}$) compared to raw PAS.

### v2

@@ 5.3 Time cost and memory usage
Generally, as depicted in the Fig. 15 and Fig. 16, the time cost of decomposition is less than 1 second in most maps and less than 3 seconds in the worst-case scenario (such as 800 to 1000 dense agents), which is deemed acceptable for practical application. In the majority of cases, all three steps contribute to the time cost. However, there are differences in the time cost of steps across different maps. In maps with an abundance of passable cells exceeding the number of agents, such as den512d, Berlin_1_256, and Paris_1_256, IC consumes almost all of the time cost, resulting in overlapping time costs after IC, BC, and LS. Conversely, in maps with few passable cells close to the number of agents, all steps contribute to the time cost, as these scenarios require BC and LS to decompose initial subproblems into smaller subproblems.
@@
5.4 Comparison with Independence Detection
Similar to Layered MAPF, Independence Detection (ID) [20] is a technique used to decompose a group of agents into the smallest possible groups. Unlike Layered MAPF, ID assigns each agent to its own unique group and then merges two groups if one group cannot find a solution (using MAPF methods) that avoids conflict with the solution of another group. ID is limited to methods that can take external paths as constraints to avoid conflicts, so not every MAPF method is applicable to ID. In this section, we compare Layered MAPF and ID using CBS and EECBS. The related results are shown in Fig. 17 and Fig. 19. For the implementation, we use the code from [30] [^3] https://github.com/svancaj/HybridMAPF for ID.
In this section, we compare Layered MAPF with ID in terms of maximum subproblem size, number of subproblems, time cost, and success rate in finding a solution. Since MAPF methods are embedded in ID, we compare the total cost of Layered MAPF (including the time cost of decomposition and solving subproblems) with the time cost of ID. To maintain consistency, we use the same MAPF instances for the decomposition of each instance.
As a common rule in solving MAPF instances, we set an upper bound on the time cost (30 seconds). Methods that are complete but fail to find a solution within the allotted time are considered to have failed. When ID fails to find collision-free solutions for an instance, we use the number of agents in the instance as its maximum subproblem size. To maintain consistency, this rule is also applied to Layered MAPF, even though our method decomposes the instance into multiple subproblems.
Figure 17: These figures show how Layered CBS and ID using CBS perform across various maps. Their performance is evaluated in terms of time cost, success rate, maximum subproblem size, and total number of subproblems. The data for Layered CBS is shown as solid lines, while the data for ID (using CBS) is shown as dotted lines. More details about the maps can be found in Fig. 15 and Fig. 16. The last row contains the legend for the upper figures; figures in the same column share the same legend, and the same applies hereinafter.
Figure 18: These figures show the average time cost, success rate, maximum subproblem size, and number of subproblems for Layered CBS and ID (using CBS) across various maps. These figures summarize the data from Fig. 17. The map indices correspond to those in Fig. 15 and Fig. 16, and the same applies hereinafter.
Figure 19: These figures show how Layered EECBS and ID using EECBS perform across various maps. Their performance is evaluated in terms of time cost, success rate, maximum subproblem size, and total number of subproblems. The data for Layered EECBS is shown as solid lines, while the data for ID (using EECBS) is shown as dotted lines. More details about the maps can be found in Fig. 15 and Fig. 16.
Figure 20: These figures show the average time cost, success rate, maximum subproblem size, and number of subproblems for Layered EECBS and ID (using EECBS) across various maps (please refer to Fig. 16 for the maps corresponding to the map indices). These figures summarize the data from Fig. 19.
As shown in Fig. 17 and Fig. 19, for both CBS and EECBS, Layered MAPF’s maximum subproblem size increases significantly slower than ID as the number of agents increases. Additionally, Layered MAPF decomposes the MAPF instance into more subproblems compared to ID. As a result, Layered MAPF has a lower time cost (for CBS, 14.5s $<$ 27.4s; for EECBS, 14.4s $<$ 26.0s), a higher success rate (for CBS, 0.64 $>$ 0.12; for EECBS, 0.65 $>$ 0.18), a smaller maximum subproblem size (for CBS, 206.8 $<$ 426.5; for EECBS, 206.9 $<$ 426.9), and more subproblems (for CBS, 218.7 $>$ 1.9; for EECBS, 218.7 $>$ 2.31), on average, compared to ID, as shown in Fig. 18 and Fig. 20.
There are several reasons that cause this phenomenon:
1, ID does not consider the order of solving subproblems, while Layered MAPF does. Therefore, Layered MAPF has a higher probability of decomposing a MAPF instance into smaller subproblems;
2, when ID tries to check whether a group (i.e., a subproblem) can avoid conflicts with another group, it does not consider avoiding groups that it has already avoided in the past. As a result, ID can sometimes fall into infinite loops. For example, suppose group A avoids group B. Then, ID detects conflicts between group A and group C and finds a new solution for group A that avoids conflicts with group C. However, this new solution for group A conflicts with group B. ID will then repeat the process of making group A avoid group B and group C until it runs out of time. This phenomenon is less frequent when there are a few sparse agents but becomes more common when there are dense agents;
3, in Layered MAPF, each agent participates in a multi-agent pathfinding process only once. In contrast, in ID, an agent may participate multiple times, as it can be merged into different groups at different stages. This causes ID to have a higher time cost than Layered MAPF.
5.5 Summary
Figure 21: This figure illustrates how the maximum subproblem size changes as the density of agents and the number of agents in the instance change. The data source is the same as in Fig. 15 and Fig. 16.
@@ 5.5 Summary
Our method demonstrates high effectiveness for MAPF instances with cells exceeding two times the number of agents, while its effectiveness diminishes as the number of passable cells approaches two times the number of agents. This is because a higher number of passable cells increases the likelihood of decomposing the instance into smaller subproblems. Conversely, for the same map, the effectiveness of decomposition decreases with an increase in the number of agents in the MAPF instance.
@@
In terms of costs, the average time cost is less than 1 second in average, with a maximum of less than 3 seconds in the worst cases (such as 800 to 1000 dense agents), and memory usage remains below 1 MB. Comparatively, maps with more passable cells than the number of agents require fewer computations and less memory space.
We also studied how the decomposition of an instance changes as the number of agents and the agent density change (defined as the ratio between the number of agents and the number of passable cells), as shown in Fig. 21. We can observe that the maximum subproblem size is large only when both the agent density and the number of agents are high. In other words, instance decomposition can be effective even with a large number of agents, as long as the agent density is low.
6 Results of decomposition’s application
Time cost (ms)
Memory Usage (MB)
Success rate
Sum of cost
Makespan
@@ 6 Results of decomposition’s application
Figure 23: These figures provide a summary comparison between the raw MAPF method and the Layered MAPF method, combining the results across all maps. In the visualizations, crossed bars represent the Layered version of the MAPF method, while empty bars represent the raw MAPF method.
[…]
In this section, experiments were conducted on a computer running Ubuntu 20.04, equipped with an Intel Xeon(R) CPU (2.1 GHz) and 64 GB of memory. All code was implemented in C++.
6.1 Explicit Estimation CBS
Explicit Estimation CBS (EECBS) [7] stands out as one of the state-of-the-art CBS-based MAPF methods, characterized by its completeness and bounded suboptimality. Further details can be found in Section 2. EECBS’s code  [^4] https://github.com/Jiaoyang-Li/EECBS operates by searching each agent’s path separately and provides interfaces to avoid conflicts with external paths, rendering it a serial MAPF method. In its Layered version, which we refer to as “Layered EECBS", previous subproblems’ paths and subsequent subproblems’ starts are set as constraints. We use the number of constraint nodes as an estimate of the memory usage of EECBS, as it is the only variable that increases during its execution.
The average time cost, success rate, memory usage, makespan, and total cost of Layered EECBS and raw EECBS are shown in Fig. 23. How Layered EECBS and raw EECBS perform under various maps as the number of agents increases is shown in Fig. A1 in the Appendix. As shown in Fig. A1, Layered EECBS has lower time cost and thus higher success rates compared to raw EECBS when there are few agents. However, these advantages slowly diminish as the number of agents increases, as does Layered EECBS’s advantage in memory usage. Despite this, Layered EECBS still has an advantage in most cases. In terms of path quality, the solutions of Layered EECBS are slightly worse than those of raw EECBS. A summary of the performance of Layered EECBS and raw EECBS can be found in Fig. 23.
As illustrated in Fig. 23, on average, Layered EECBS exhibits lower time costs ($1.2\times 10^{4}$ ms $<$ $1.7\times 10^{4}$ ms), leading to a higher success rate (0.68 $>$ 0.57). The time cost of decomposition is relatively small compared to the time cost of solving the MAPF instance. Regarding memory usage, Layered EECBS consumes less memory space compared to raw EECBS (4839.0 constraint nodes $<$ 7542.8 constraint nodes). Decomposition of the MAPF instance significantly reduces the growth of the constraint tree’s size by decreasing the number of agents solved simultaneously.
In terms of path quality, Layered EECBS yields a larger makespan (Layered EECBS: $2.3\times 10^{2}$, raw EECBS: $1.9\times 10^{2}$), resulting in a larger sum of costs (Layered EECBS: $4.7\times 10^{4}$, raw EECBS: $3.2\times 10^{4}$). In maps where both raw EECBS and Layered EECBS achieve a success rate close to 1 (e.g., map 1, 8, 10, 19, 21 and 22), both methods exhibit similar makespan and sum of costs. However, in maps where raw EECBS has a lower success rate compared to Layered EECBS (e.g., map 4, 16, 17 and 24), Layered EECBS demonstrates a larger makespan and sum of costs, as depicted in Fig. 22.
@@ 6.1 Explicit Estimation CBS
6.2 Priority Based Search
Priority Based Search (PBS) [4] is an incomplete, suboptimal, and priority-based MAPF method. PBS searches each agent’s path separately and it’s code [^5] https://github.com/Jiaoyang-Li/PBS provides interfaces to avoid conflicts with external paths, making it a serial MAPF method. In its Layered version, which we refer to as “Layered PBS", we set previous subproblems’ paths and subsequent subproblems’ starts as constraints. We use the number of priority nodes as an estimate of the memory usage of PBS, as it is the only variable that increases during its execution.
The average time cost, success rate, memory usage, makespan, and total cost of Layered PBS and raw PBS are shown in Fig. 23. How Layered PBS and raw PBS perform under various maps as the number of agents increases is shown in Fig. A2 in the Appendix. As shown in Fig. A2, Layered PBS has lower time cost and thus higher success rates compared to raw PBS when there are few agents. However, these advantages slowly diminish as the number of agents increases, as does Layered PBS’s advantage in memory usage. Despite this, Layered PBS still has an advantage in most cases. In terms of path quality, the solutions of Layered PBS are slightly worse than those of raw PBS. A summary of the performance of Layered PBS and raw PBS can be found in Fig. 23.
As shown in Fig. 23, on average, Layered PBS has a lower time cost ($1.3\times 10^{4}$ ms $<$ $2.0\times 10^{4}$ ms) and thus a higher success rate (0.68 $>$ 0.37). The time cost of decomposition is relatively small compared to the time cost of solving the MAPF instance. In terms of memory usage, Layered PBS has lower memory usage compared to raw PBS (234.4 priority nodes $<$ 706.8 priority nodes) because decomposition of the MAPF instance reduces the complexity of the priority tree significantly.
When it comes to path quality, Layered PBS has a larger makespan ($2.3\times 10^{2}$ $>$ $1.0\times 10^{2}$) and thus a larger SOC ($4.5\times 10^{4}$ $>$ $1.3\times 10^{4}$). Similar to EECBS, raw PBS finds it difficult to find longer solutions and can only finds shorter solutions, resulting in lower success rates for maps where raw PBS struggles (e.g., map 4, 6, 16 and 17). For maps where both methods have high success rates (e.g., map 1, 9 and 10), they have similar makespan and SOC, as shown in Fig. 22.
@@ 6.2 Priority Based Search
6.3 Large Neighborhood Search 2
Large Neighborhood Search 2 (LNS2) [3] is a suboptimal MAPF method that starts from a set of paths with collisions and repeatedly replans subsets of paths to reduce the overall number of collisions until the paths become collision-free. LNS2 updates each agent’s path separately, and its code [^6] https://github.com/Jiaoyang-Li/MAPF-LNS2 provides interfaces to avoid conflicts with external paths, making it a serial MAPF method. In LNS2’s implementation, LNS2 considers all agents in the initial planning phase, so it generally has the highest memory usage during this phase. LNS2 provides several classic MAPF methods to generate the initial solution, including EECBS, PP, PPS, CBS, PIBT, and winPIBT. To obtain a conflict-free solution as quickly as possible, we choose EECBS to generate the initial solutions. Therefore, we use the number of constraint nodes in EECBS during the initial planning phase as an estimate of the memory usage of LNS2, as this is the variable that contributes most to memory usage during its execution.
The average time cost, success rate, memory usage, makespan, and total cost of Layered LNS2 and raw LNS2 are shown in Fig. 23. How Layered LNS2 and raw LNS2 perform under various maps as the number of agents increases is shown in Fig. A3 in the Appendix. As shown in Fig. A3, Layered LNS2 has lower time cost and thus a higher success rate when there are few agents compared to raw LNS2. However, these advantages slowly diminish as the number of agents increases, as does Layered LNS2’s advantage in memory usage. Despite this, Layered LNS2 still has an advantage in most cases. In terms of path quality, the solutions from Layered LNS2 are slightly worse than those from raw LNS2. A summary of the performance of Layered LNS2 and raw LNS2 can be found in Fig. 23.
As shown in Fig. 23, on average, Layered LNS2 has a lower time cost ($1.2\times 10^{4}$ ms $<$ $1.6\times 10^{4}$ ms) and thus a higher success rate (0.72 $>$ 0.52). In terms of memory usage, Layered LNS2 has lower memory usage compared to raw LNS2 (24.9 constraint nodes $<$ 228.8 constraint nodes). Due to the decomposition of the MAPF instance, other agents’ paths are considered as dynamic obstacles and do not join conflict resolve, resulting in Layered LNS2 requiring fewer repair actions to fix conflicts between agents’ paths.
Regarding path quality, Layered LNS2 has a larger makespan ($2.4\times 10^{2}$ $>$ $1.9\times 10^{2}$) and SOC ($4.7\times 10^{4}$ $>$ $3.0\times 10^{4}$). Similar to EECBS and PBS, raw LNS2 struggles to find longer solutions and can only discovers shorter solutions, particularly in maps where raw LNS2 has a lower success rate (e.g., map 12 and 16). For maps where both methods have similar success rates (e.g., map 1, 2, 8, 21 and 22), raw LNS2 and Layered LNS2 have similar makespan and SOC, as shown in Fig. 22.
@@ 6.3 Large Neighborhood Search 2
6.4 Push and Swap
Push and Swap (PAS) [31] is a suboptimal and complete rule-based MAPF method. Push and Swap updates agents’ paths simultaneously (via an action called “MULTIPUSH”, which moves a set of adjacent agents simultaneously), and its code [^7] https://github.com/Kei18/pibt2 provides no interfaces to avoid conflicts with external paths and ensures that the same state of agents does not occur twice in the solution, making it a parallel MAPF method. In its Layered version, referred to as “Layered Push and Swap" or “Layered PAS," we set previous subproblems’ targets and subsequent subproblems’ starts as static obstacles. Subsequently, we merge subproblems’ solutions by adding wait actions after solving all subproblems. In the implementation of Push and Swap, the variable that occupies the most memory is a two-dimensional matrix that stores each agent’s location at all times. Therefore, we use the number of elements in the matrix as an estimate of the memory usage of Push and Swap, which is calculated as the makespan of the solution multiplied by the number of agents.
The average time cost, success rate, memory usage, makespan, and sum of cost of Layered Push and Swap (PAS) and raw PAS are shown in Fig. 23. How Layered PAS and raw PAS perform under various maps as the number of agents increases is shown in Fig. A4 in the Appendix.
As shown in Fig. A4, in some cases, Layered PAS takes more time than raw PAS, while in other cases, it does not, as PAS’s time cost is close to the time cost of instance decomposition. However, when there are many agents, Layered PAS takes less time than raw PAS. Consequently, Layered PAS achieves a higher success rate than raw PAS as the number of agents increases in most cases.
In terms of memory usage, Layered PAS consistently requires less memory than raw PAS, even as the number of agents increases.
Regarding path quality, Layered PAS’s solutions are slightly worse than raw PAS’s solutions. A summary of Layered PAS and raw PAS’s performance can be found in Fig. 23.
As shown in Fig. 23, on average, Layered PAS consumes less time ($6.8\times 10^{3}$ ms $<$ $1.15\times 10^{4}$ ms) and memory space ($6.0\times 10^{6}$ constraint nodes $<$ $4.0\times 10^{7}$ constraint nodes) than raw PAS, resulting in a higher success rate (0.92 $>$ 0.69). Due to the number of “push” and “swap” actions grows exponentially as the number of agents increases, decomposing the MAPF instance enables Layered Push and Swap to solve subproblems separately, requiring fewer “push” and “swap” actions.
Regarding path quality, as PAS is a parallel MAPF method, the addition of wait actions to merge subproblems’ solutions, Layered PAS has a larger makespan ($1.0\times 10^{4}$ $>$ $4.0\times 10^{3}$) and SOC ($3.5\times 10^{6}$ $>$ $1.2\times 10^{6}$) compared to raw PAS.